OpenStudy (anonymous):

***CHECK*** What is the exact volume of a cylinder that has a radius of 9.2 in. and a height of 3.7 in.? ___pi in3

OpenStudy (anonymous):

I think its 983.34

OpenStudy (fibonaccichick666):

Can you tell me what you did

OpenStudy (fibonaccichick666):

cause that's a bit high

OpenStudy (fibonaccichick666):

(did you maybe not leave your answer in terms of pi?)

OpenStudy (anonymous):

Yea I used the formula \[v=\pi r^2h\]

OpenStudy (anonymous):

\[\pi \times 9.2^2\times3.7\]

OpenStudy (fibonaccichick666):

so, you don't want to multiply by pi

OpenStudy (fibonaccichick666):

because it asks for the exact answer, it wants it in terms of something times pi

OpenStudy (fibonaccichick666):

that's why the answer input has \(\pi in^3\)

OpenStudy (anonymous):

Oh I did multiply pi in my equation\[\pi \times 9.2^2\times3.7\]

OpenStudy (fibonaccichick666):

yea, it wants just the others multiplied together, not with pi

OpenStudy (anonymous):

Oh ok i get it now

OpenStudy (fibonaccichick666):

something like \(32\pi in^3 \)would be the style of the answer

OpenStudy (fibonaccichick666):

So, what do you get for yours?

OpenStudy (anonymous):

Ok I didnt multiply pi and I got 313.168

OpenStudy (fibonaccichick666):

that sounds plausible

OpenStudy (anonymous):

Ok thx c:

OpenStudy (fibonaccichick666):

and my calculator agrees Np